Lenovo uses the ZX-81 chip in some of its laptop computers. The prices for the chip during the last 12 months were as followsneed help with the 3rd part
Answer #1

Where F(t-1) is the forecast for the previous period and A(t-1) is the actual demand for the previous period.

Forecast for November= Forecast for october+ alpha(actual price for october-forecast for October)

= 1.73 + (0.5* (1.57 - 1.73)) = 1.65
Forecast for December = Forecast for December+ alpha(actual price for November-forecast for November)

= 1.65+ (0.5 * (1.62 - 1.65)) = 1.63